Tracing of curve y^2=x(x-2)^2
1
Expert's answer
2020-05-05T20:21:04-0400

Given curve "y^2 = x(x-2)^2".

At "x=0 \hspace{0.05 in} and \hspace{0.05 in} x= 2, y = 0."

Also "y^2 \geq 0" for all value of x.

Given curve is symmetric about x-axis as by putting y = -y curve remain unchanged.

So, let discuss only positive side of y-axis. Then Negative side of y-axis plot symmetrically around x-axis.

So "y = \sqrt{x} (x-2)".

Now, Differentiate given curve with respect to x, we get

"y' = \sqrt{x} + \frac{x-2}{2\sqrt{x}}"

So, "y' > 0" for "x > 2". Hence, y is an increasing function after x=2.

Now "y' = 0 \implies 2x+x-2 = 0 \implies x = 2/3 = 0.66."

And "y'>0" for "0 \leq x < 0.66" , "y' <0" for "0.66<x<2".